2A + 3B --> 4C,

20.0 g of A reacted with excess of B. How many grams of C were obtained, if the molar mass of A is 26.0 g/mol and molar mass of C is 14.5 g/mol?


1
Expert's answer
2021-10-20T02:52:08-0400

n(A) "= \frac{20.0}{26.0} = 0.7692 \;mol"

According to the reaction:

n(C) = 2n(A) "= 2 \times 0.7692 = 1.5384 \;mol"

m(C) "= 1.5384 \times 14.5 = 22.30 \;g"

Answer: 22.30 g